### v1

# Matthias Schonlau

Abstract
Effective methods for visualizing data involving multiple variables, including categorical ones, are limited. The hammock plot (Schonlau, 2003) visualizes both categorical and numerical variables using parallel coordinates. We introduce the Stata implementation hammock. We give numerous examples that explore highlighting, missing values, putting axes on the same scale, and tracing an observation across variables. Further, we introduce parallel univariate plots as an edge case of hammock plots. We also present and make publicly available a new dataset on the 2020 Tour de France. A graphical abstract is shown below.
keywords:
st0001, visualization, hammock plot, clustergram, parallel coordinate plot, alluvial plot, Sankey diagram
1 Introduction
Leonardo da Vinci wrote, a poet would struggle to “describe with words what a painter is able to [depict] in an instant” (Janson and Janson, 2004). But not everybody can paint like da Vinci. Likewise, it is often easier to write a few pages of text than to construct a sensible graph. Arguably, readers appreciate graphs more than text. Empirical evidence suggests, graphs even lend an aura of scientific credibility (Tal, 2015). For those of us who cannot paint like da Vinci, software makes it easier to create graphs. In this paper we introduce software for creating hammock plots.
For visualizing one or two variables there are many choices: bar charts, stacked and grouped bar charts, scatter plots, pie charts and others. A scatter plot matrix and a parallel coordinate plot are effective for visualizing more than two continuous variables. Mosaic plots (Hofmann, 2008) are suitable for more than two categorical variables, though they are harder to interpret. Another option is to show multiple plots that are linked through highlighting. For displaying multiple variables that contain a mixture of categorical or numerical variables in a single plot, we need a hammock plot or one of its variations. This paper showcases the hammock plot and its Stata implementation.
To introduce a less familiar plot, it may be easiest to give an example using a data set that many Stata users are familiar with, the auto data. In this data set, each of the 74 observations represent one car. Each car’s repair record was rated with 1-5 stars, 5 being best. Figure 1 shows two plots:
Figure 1: A scatter plot and a hammock plot of repair record (1-5 “stars”) vs car origin from the auto data. With only two axes, the usefulness of the hammock plot may not yet be apparent.
a scatter plot of the variables repair record and car origin (foreign vs domestic) and a corresponding hammock plot. The repair record takes the values 1 through 5 corresponding to 1 through 5 “stars”. To reduce overplotting, the scatter plot uses jittering, i.e., we add a little bit of random noise. A scatter plot is not the most effective way to display these two variables; our interest is in the layout of the axes. In a scatter plot the two axes are aligned at a right angle. A hammock plot uses parallel coordinates: the axis for repair record is parallel to the axis for car origin. On the axes, there are bars with univariate statistics; the height of the bar is proportional to the number of observations it contains. Axes are connected by bars. The width of a connecting bar is also proportional to the number of observations a bar contains.
A hammock plot with two axes may not be particularly useful. However, parallel axes have the advantage that we can add additional axes. Figure 2 shows a hammock plot (Schonlau, 2003) of a half dozen variables of the auto data set. (We chose some variables to illustrate the hammock plot; a different set of variables could have been chosen. The Stata code for this plot is given later in Section 3.1.)
Figure 2: Hammock plot of the auto data. Foreign cars are highlighted in orange. A category for missing values, where present, is displayed near the bottom. Here, the dashed black line was added to show the separation between the areas for missing and non-missing values.
The graph consists of columns with univariate information and lines/ bars that connect the columns. We discuss each in turn.
Each variable is shown in a column or axis; the variable name is displayed beneath the column. For categorical variables, each column contains a bar that corresponds to one category. The length of the bar is proportional to the number of observations it represents. For example, the variable foreign in the last column has two categories, “foreign” and “domestic” representing foreign and domestic cars. We see there are more than twice as many domestic cars as foreign cars in this data set. For numerical variables, there is also a “bar” for each value. However, for values occurring only once, the bar is so thin that it looks like a line. So, the univariate display for a numerical value looks like a rug plot (see the discussion in Cox (2025)), which we typically see at the bottom of a density plots to show the individual data points of a continuous variable.
“Missing values” is a separate category and, optionally, a space for missing values is reserved at the bottom. Figure 2 features an optional line that separates the area between missing and non-missing values. The variable repair record has a bar at the bottom which indicates its missing values.
Neighboring axes are connected by bars. In Figure 2, the connecting bars are rectangles, but other shapes could be used (e.g., parallelograms, which are faster to compute). The width of the rectangles is proportional to the number of observations it contains. Rectangles containing very few observations appear as a line. The rectangles each represent a cell of the two-way contingency table of the two adjacent variables, i.e., they represent bivariate statistics.
Figure 2 highlights foreign cars in orange. We can follow the observations belonging to foreign cars throughout the graph. For example, we see that foreign cars tend to have less headroom, a better repair record, lower weight, and higher mileage. For the continuous variables mileage (miles per gallon) and weight we can also make a statement about correlations. Mileage and weight are negatively correlated because the lines cross in opposite directions: high values of mileage correspond to low values of weight and conversely. In fact, the correlation between mileage and weight turns out to be 0.80.8-0.8- 0.8.
The remainder of this paper is as follows. Section 2 gives a brief history of the hammock plot and similar plots that followed. Section 3 gives numerous examples and explains the Stata syntax in the context of the examples. Section 4 concludes.
2 A brief history of the hammock plot and related plots
In 2003, Schonlau (2003) proposed the hammock plot to visualize both categorical and numerical variables. The hammock plot was inspired by the clustergram (Schonlau, 2002, 2004), a program to visualize how clusters form as you increase the number of clusters. Like the parallel coordinate plot (Inselberg, 1985; Wegman, 1990) for continuous variables, the hammock plot also uses parallel axes. However, unlike parallel coordinate plots, hammock plots use two-dimensional plotting elements rather than lines to connect adjacent axes. A hammock plot of only continuous variables (without labels, univariate statistics, and missing values) is a parallel coordinate plot. More recently, Schonlau (2024) proposed connecting neighboring axes with rectangles rather than parallelograms. This avoids the so-called reverse line-width illusion.
Several hammock-type plots have been proposed including the ParSets plot (Kosara et al., 2006), alluvial plot (Rosvall and Bergstrom, 2010), common angle plot (Hofmann and Vendettuoli, 2013), and the cpcp plot (Pilhoefer and Unwin, 2013). Such plots differ in how they connect two adjacent categorical variables (e.g., rectangle vs. rounded shapes), two adjacent numerical variables (e.g., rectangle vs. line), two adjacent numerical-categorical variables (e.g., rectangle vs. triangle), whether they allow numerical variables or force categorical encodings, and how they space out numerical variables along a numerical axis.
The generalized parallel coordinate plot (GPCP) (VanderPlas et al., 2023) uses individual lines (rather than areas) to connect neighboring axes. The plot resolves ties by plotting parallel lines (rather than areas). For small data sets, this allows tracing individual observations across the plot. To trace individual observations, the hammock plot uses highlighting (see Section 3.2).
Perhaps the best known plot among the hammock-type plots is the alluvial plot (Rosvall and Bergstrom, 2010). Whereas the hammock plot uses bars (parallelograms or rectangles), alluvial plots use rounded curves for the bivariate elements that connect axes. These plots were initially only used to visualize network variables over time. Alluvial plots are now also used to visualize categorical variables; they do not accommodate numerical variables. In Stata, alluvial plots are implemented in the package alluvial Naqvi [2024a].
A similar looking plot, the Sankey diagram (Sankey, 1898; Schmidt, 2008), visualizes flows of energy and materials rather than variables. A historical Sankey diagram is shown in Figure 3.
Figure 3: A Sankey diagram of flows of energy (Sankey, 1898)
A Sankey diagram may have circular flows or one flow may stop in the middle of the graph. The Sankey diagram is implemented in Stata (Naqvi, 2024b). The names “Sankey diagram” and “Alluvial plot” are sometimes conflated in the literature. If a plot visualizes variables, it is not a Sankey diagram. A more detailed overview over all of these plots can be found in Schonlau (2024). A Stata implementation of the parallel coordinate plot is in parplot (Cox, 2007).
3 Examples and Syntax
We will introduce the syntax by example. Our first example features a mixture of numerical and categorical variables, highlighting, and missing values (Section 3.1). In the following section, Section 3.2, we trace individual observations in a dataset related to the Tour de France. The variables in this example all use the same scale (seconds). The samescale option shows the effect of standardizing the layout across axes. Section 3.3, contains an example where our interest lies in connecting bars between axes; the univariate bars are largely redundant and are omitted. Section 3.4 shows the other extreme: parallel univariate bars without connectors. The examples include a graphical representation of socio-demographic table, enriching or replacing the typical “Table 1” in empirical studies.
3.1 Highlighting
The hammock plot in Figure 2 highlighted foreign cars. It was created with the following command:
  • . sysuse auto, clear
    (1978 automobile data)
    . label var rep78 "repair record"
    . label var mpg mileage
    . label var weight weight
    . label var headroom headroom
    . hammock mpg weight foreign rep78  headroom price, ///
    >  missing ///
    > hivar(foreign) hival(1) ///
    > label_format(%8.0g) ///
    > yline(10)
The option missing adds room at the bottom across all variables for a category for any missing values. We highlight the variable and specific values of that variable with hivar and hival. The label on the plot is not the number highlighted (1), but its label (“foreign”). The option label adds the labels and the univariate bars. Likewise, the variable name shown below each axis is the variable label. So, we see the label “mileage” rather than the variable name “mpg”. We can highlight (hival(0)) a single value, missing values (hival(.)), a range (hival(>27)) or several values in different colors (e.g., hival(1 4 5)). It is also possible to specify an external variable for highlighting, i.e., a variable that is not among the variables in the list of variables shown in the graph.
By default, the labels of the variable price in Figure 2 are written in scientific format if the numbers get too large. We can change this by adjusting the format in the label_format option.
Specifying Colors
Figure 2 relied on the default colors. We can specify different colors in the option colorlist. Figure 4) shows the auto data where cars with a large gas mileage are highlighted (specifically, mileage>27). We see cars with high mileage are cheaper, they tend to have lower weight, a better repair record (rep78), and less headroom.
Figure 4: Hammock plot of the auto data. Cars with high mileage are highlighted. The text explains how to choose colors rather than rely on default colors.
Figure 4 was created with the following command:
  • . hammock mpg weight foreign rep78  headroom price, ///
    > missing ///
    > hivar(mpg) hival(>27) ///
    > label_format(%8.0g) ///
    > colorlist(gs11 magenta )
The first color in option colorlist is the default color; any subsequence colors are used for highlighting in the order specified in hival(). You can list all named colors in Stata with graph query colorstyle; it is also possible to specify colors in RGB format.
We can also include colors created from the palette command. The palette command saves colors in RGB format (e.g., ”0 0 255”) in r(p) . We can input this directly into the colorlist option
. colorpalette lin vegetables
. hammock x1 x2 x3, hivar(type) hival(1 2 3) colorlist(‘" ‘=r(p)’ "’)
3.2 Tracing individual observations
Highlighting can also be used to trace individual observations across variables. Figure 5 shows the finishing time for the first and the last several stages of the 176 riders of the 2020 Tour de France (see Appendix A for details). The Tour de France is a bicycle race — arguably the best-known bicycle race — and is completed in 21 stages. Each stage corresponds to one racing day. Riders that finish as a group (without a visible gap of 1 second or at least one bike length) receive the same finishing time. At the very bottom of Figure 5 for each axis, there are missing values (riders who quit the race). Near the bottom, above the missing values, is the fastest stage time for that day, and at the top the slowest stage time for that day.
The main group of riders is called the peloton. In bicycling, it is advantageous to ride in groups to draft off other riders. You can see the peloton across the stages. Additional groups may form as riders try to break out ahead of the peloton (break-away groups and chase-groups) or cannot keep up (the so-called “autobus” group). In particular, in the mountain stages several groups may form. Mountain stages are those with variable names ending in an “m”.
Stage 20 is a time trial indicated by the variable name ending with a “t”. In a time trial, riders ride individually and cannot draft off their teammates. There are no groups in the time trial.
Looking at the riders with “missing” finishing times, i.e., riders who had to quit the race for one reason or another, we find a surprise: In stage 17, there is a line leading out of the missing group, suggesting a rider did not finish stage 17 but re-entered the race in stage 18. This rider is Bryan Coquard and upon investigation it turned out that this is an error at the data source. We did not correct this here to showcase the plot’s ability to find such errors.
We can trace the eventual winner, Pogacar, by highlighting his stage times. Pogacar was not always in the leading group. In stage 16 and 19 he let a number of riders pass. Pogacar did have the fastest time trial in stage 20 though.
Figure 5: Hammock plot of the first and the last stages of the tour de France data. The eventual winner, Pogacar, is highlighted. This plot scales each axis individually from fastest to slowest rider. Times are not comparable across axes.
Figure 5 was created as follows:
  • .         hammock s1  s15 s16 s17 s18 s19 s20 s21, ///
    >         missing ///
    >         nolabel ///
    >         colorlist(gs10 red blue yellow)   ///
    >         barwidth(.2) minbarfreq(5) ///
    >         uni_fraction(.15) ///
    >         text(97 1 "slow" 10 1 "fast"  3 1 "missing") ///
    >         hivar(rider) hival("Pogacar")
We removed the labels (nolabel) to avoid clutter. barwidth(.2) reduces the width of the bars connecting the axes (relative to the default barwidth(1)). Sometimes connecting bars of individual observations are barely visible because the bar is too thin. minbarfreq(5) increases the width of any connecting bars with fewer than 5 observations as if they had 5 observations. That is, bars with 5 or fewer observations all have the same width. Also note the variable used for highlighting, rider, is not among the variables plotted.
Figure 5 nicely shows fast and slow riders. However, a unit of time is shown differently on each parallel axis: the difference between the slowest time (at the top) and the fastest time (at the bottom, above the missing values) is not constant across stages. We subtract the time of the fastest rider for each stage and denote the new variables d1-d21. These new variables represent how many seconds have passed since the stage winner crossed the finishing line. We can then use the option samescale(varlist) to enforce the same scale for all variables. The revised plot is shown in Figure 6.
Figure 6: Hammock plot of time lost to stage winners for multiple stages. The higher a rider on the plot, the more time he lost. All axes are on the same scale. The eventual winner, Pogacar, is highlighted.
We see that riders lost a lot more time in stages 15-18 relative to the stage winner. There was little action on day 19: most riders finished in the peloton. On the last day, day 21, it is customary for all riders to ride as a group except for the sprint at the end. This is why stage 21 shows one big peloton with other riders a few seconds behind.
For the revised plot, Figure 6 , we first computed the d1-d21 variables by subtracting the fastest stage time. The variable name is changed from s‘i’ to ‘i’ with d‘i’ ranging from 1 to 21. THe code is as follows:
  • .         foreach var of varlist s* {
      2.                 quietly summarize `var´, meanonly
      3.                 local newname = "d" + substr("`var´", 2, .)
      4.                 quietly gen `newname´ = `var´ - r(min)
      5.         }
    .         hammock d1 d15 d16 d17 d18 d19 d20 d21, ///
    >         nolabel ///
    >         missing ///
    >         colorlist(gs10 red blue yellow)   ///
    >         barwidth(.2) minbarfreq(5) ///
    >         uni_fraction(.15) ///
    >         hivar(rider) hival("Pogacar") ///
    >         samescale(d1 d15 d16 d17 d18 d19 d20 d21)
3.3 A connectors-only Hammock plot
Sometimes it makes sense to remove the univariate charts and concentrate on the connectors. Figure 7 shows marital status over 5 time points (or waves) in the GESIS panel (https://www.gesis.org/en/gesis-panel), a survey panel in Germany. A survey panel is a longitudinal survey where the same individuals (or units) are surveyed repeatedly over time. The most common category in Figure 7 is “married” (living together rather than apart), followed by “single”. We can see that marital status does not change much within a few short years. What might surprise is the small subset of observations that went from “married” to “single” (rather than “divorced” or one of the other categories). It turns out that marital status is self-assessed and some respondent preferred “single” over “divorced” or another category.
Figure 7: Marital status over 5 waves (or time points) in the GESIS panel. Surprisingly, some married individuals return to “single” rather than to “divorced” (or similar) the following year.
For comparison, Figure 8 shows a plot that includes the univariate bars. Figure 7 was created as follows:
  • . hammock wave1 wave2 wave3 wave4 wave5, ///
    > nolabel ///
    > nounibar ///
    > barwidth(3) ///
    > space(0.1) ///
    > color(blue%30) ///
    > text(16 1 "single" ///
    >      38 1 "married " ///
    >      57 1 "married apart" ///
    >      80 1 "divorced" ///
    >     100 1 "widowed" ) ///
    > graphregion(margin(l+10 r+10))
Figure 8: Comparison plot to Figure 7 with univariate bars included.
nounibar requests not to plot bars on the axes. space(0.1) leaves some space between the connectors. We specified nolabel because the same label applies for all variables. Instead, we added the labels by hand using Stata’s text option which is passed through to the plotting command. The x-coordinate for the text is always $11$ because the text is for the first x-variable and the first x-variable has the x-coordinate $11$. The y-coordinates are spaced between 0 and 100; the exact values are determined by trial and error. (Temporary optional arguments such as yline(10 20 30 40 50) can help finding desired coordinates). barwidth(3) triples the width of the connecting bars. We add some space on the left and right using Stata’s graphregion(margin()) option. This avoids that some text runs off the page. For variety, we specified a different color using a color intensity of 30%. We did not reserve an area for missing values in this plot.
More on the size and shape of bars
uni_fraction(0.5) controls the length of the univariate bars. By default, 50% of the vertical space is filled with univariate bars. If any two bars overlap, this fraction can be adjusted downward to avoid the overlap. This fraction can also be adjusted upward to suit. Likewise, barwidth(1) can be adjusted to control the width of the connectors to reduce visual clutter or to reveal more detail. The number specified is an adjustment factor relative to the default, 1 (see an example in Figure 7). Connecting bars that represent a very small fraction of observations can be so thin that they are barely visible or do not render properly. As seen in the code for Figure 5, option minbarfreq(1) specifies that a rectangle must represent at least minbarfreq observations. All bars corresponding to fewer than minbarfreq observations are displayed as if they contained minbarfreq observations. The default value, 1, does nothing because all rectangles contain at least 1 observation.
The option shape(rectangle) controls the shape of the connectors: rectangle or parallelogram. The main advantage of the parallelogram is computational speed. The rectangle shape fixes the so-called reverse line width illusion (see Schonlau (2024)). The helpfile describes additional layout options.
3.4 Parallel univariate plots
A hammock plot shows an alternating sequence of univariate plots (univariate bars on the axes and bivariate plots (connecting bars between two neighboring axes). If we omit the bivariate plots, parallel univariate plots emerge as an edge case. In Stata, we simply specify space(1). Figure 9 shows an example for the 2020 Tour de France data. Without the connecting bars we are able to display more variables. We see that the winner, Pogacar, was not always in the winning group, but except for stage 16 he was never far behind.
Figure 9: Parallel univariate plot of the 2025 Tour de France. The winner of the tour, Pogacar, is highlighted. Except in stage 16, Pogacar was never far behind the stage winners.
The parallel univariate plots are connected with one another through highlighting. The plot was constructed with the following command:
  • .         hammock d*, ///
    >         nolabel ///
    >         colorlist(gs10 red blue yellow)   ///
    >         barwidth(.2) minbarfreq(5) ///
    >         uni_fraction(.15) ///
    >         hivar(rider) hival("Pogacar") ///
    >         samescale(d*) ///
    >         space(1)
In empirical studies, “Table 1” often shows the demographic composition of the study participants. Such a “Table 1” is shown in Table 1 with percentages for categorical variables and mean/ standard error for continuous ones. The data are from an asthma study (Schonlau et al., 2005). Alternatively, one might show a parallel univariate chart plot as shown in Figure 10. We see a graphical representation of separate demographic x-variables. The univariate charts not linked in any way. However, we could easily highlight a specific category, such as female participants, and observe the percentage of female participants in each category.
Variable Category Percent
Race black 6.5
hispanic 14.6
white 69.2
other race 9.7
Income <<< 15K 47.2
15K $\leq$ income<<< 30K 26.7
30K $\leq$ income<<< 50K 15
50K $\leq$ income 11.1
Education <=<=< = 8th grade 9.2
some high school 17.8
high school 29.2
some college 32.4
4yr college 6.0
>=>=> = 4yr college 5.4
Gender male 20.0
female 80.0
Age mean 42.8
SE 16.8
Severity intermittent 40.5
mild 23.8
moderate 21.1
severe 14.6
Comorbidities mean 1.57
SE 1.38
Insurance no 90.8
yes 9.2
Table 1: Demographic information about the study population in a study related to asthma. The same data are shown in parallel univariate plots in Figure 10.
Figure 10: Parallel univariate plot of the asthma data. The same data are shown as a table in Table 1.
The asthma plot corresponding to Table 1 was constructed with the following command:
  • . hammock race income education male age severity comorb no_ins, ///
    >   space(1) ///
    >   uni_fraction(.36) ///
    >   labelopt(size(vsmall)) xlab(,labsize(small))
The value for option uni__\__fraction (which affects the relative size of the bars) was chosen as large as possible without the two bars corresponding to 0 and 1 comorbidities touching. (If we move a bar upwards, then the variable would become categorical because the meaning of distance is lost.)
3.5 Re-ordering variables and categories
The variable order can be rearranged to make the overall plot more understandable or to make a specific point. As a starting point, one may arrange the order to maximize the correlation of neighboring variables. For the auto data, Figure 2 shows a variable ordering starting with mileage (on the left), where the following variable is chosen to maximize the correlation with the current variable. For example, among the remaining variables, weight has the highest correlation with mileage, “car origin” has the highest correlation with weight, and so forth.
Figure 2 shows mileage and weight are negatively correlated; domestic cars are heavier than foreign cars, and the repair record of foreign cars is better. The relationship with price is not obvious, and we might start to highlight expensive cars.
To better understand why the ordering in Figure 2 may be a sensible starting point, we now do the opposite: again starting with mileage, we continue with the variable with the lowest squared correlation to mileage, and so forth. The resulting plot is in Figure 11.
Figure 11: Hammock plot of the auto data: variables are ordered to minimize the squared correlation among neighboring variables. Instead we prefer the ordering in Figure 2.
We believe that this Figure is somewhat harder to read. The relationship among the first five variables is less obvious. We can still see that foreign cars tend to weigh less as these two variables remain in the same order.
Re-arranging categories
When the variable is continuous or ordered categorical, the order of the values and categories is fixed (in rare circumstances it may be sensible to reverse the order). When a variable is unordered categorical, we can re-arrange the categories to make the graph more readable or to bring out an aspect of the story we wish to tell. Often, a graph is easier to read when — among neighboring variables — low bars connect to low bars and high bars connect to high bars.
The only unordered categorical variable in Figure 2 is car origin (foreign vs domestic). In the current order, connecting bars from domestic (at the bottom) connect to low repair bars (at the bottom). This accentuates the finding that domestic cars have worse repair records.
4 Conclusion
The hammock plot visualizes categorical and numerical variables in the same plot. The plot is often most effective when a category or variable range is highlighted. Among the variations of the hammock plot is the alluvial plot, developed later and unable to accommodate numerical variables. The hammock plot or one of its variations is not the only way to gain insight into data. But it is one more tool in your toolbox and it is often useful.
The parallel univariate plots in Figures 9 and 10 displayed bars. Future work may extend such parallel univariate plots by displaying a stacked bar chart for each x-variable instead of the bars, or a boxplot/ violinplot for continuous variables such as age. Parallel univariate plots that do not use highlighting can be constructed from marginal statistics alone (e.g., Figure 10). Parallel univariate plots that are linked through highlighting such as Figure 9 require bivariate statistics (e.g., variables “rider” and the stage time “s1” for stage 1). Because we usually want to experiment with highlighting more than one variable, parallel univariate plots that use highlighting effectively require the whole data set.
I (Schonlau) have developed the hammock command since 2002. For the most recent version see the Boston SSC archive (ssc install hammock) or my github page (Schonlau, 2025).
Appendix A: 2020 Tour de France data
The 2020 Tour de France data is made available as part of this paper. They were scraped from the website https://www.bikeraceinfo.com/tdf/tdf2020.html by the authors. The data contain rider name, team name, 21 variables (s1-s21) for the stage finishing times, and 21 variables for the cumulative finishing times (c1-c21). All finishing times are given in seconds. Variable labels (not: variable names) of mountain stages and time trials are denoted with “m” and “t”, respectively.
Cumulative times were also scraped from the website listed above. They differ from the sum of the individual times: for the 2020 Tour de France, the first/second/third place winner of any stage was awarded a reduction of 10/6/4 seconds to their overall time.
As pointed out earlier, the entry for Stage 17 for Bryan Coquard is missing. Based on the website https://www.procyclingstats.com, he finished 35min 45sec after the stage winner. His value for s17 should be 2145 (seconds).